OpenStudy (anonymous):

The monthly payment for a car lease is $530 for 18 months. The down payment is $2,000. The residual value is $7,000. What is the total cost of the lease? Select one of the options below as your answer: A. $18,540 B. $20,540 C. $25,540

OpenStudy (anonymous):

530 * 18 = 9540 9540 + 2000 = 11540 11540 + 7000 = 18540

OpenStudy (anonymous):

A is the answer